5.a) Given a vector space V, when is a subset T ⊆ V of V said to be a vector subspace of V? 


b) Given a linear transformation V "\to" W show that the kernel Ker f of f is a subspace of V


c) Given any linear transformation

V "\to" W show that the subsetI m[f] ={y∈W y=f(x), for some x∈V} is a vector subspace of W


d) Given any linear transformation V "\to" W, show that f is one-to-one, if and only if, its kernel is trivial i.e, Ker f={0}


e) Given a linear transformation V "\to" W between finite dimensional vector spaces, show that dim Ker f + dim Im[f] =dim V

Hint: Show Ker f is finite dimensional and then starting with a basis Bo for Ker f extend to a basis of V


1
Expert's answer
2021-10-27T06:29:19-0400

(a) A subset T of a vector space V is a subspace of v if:


- T is non-empty


-For every v,t "\in" T and scalars

a,b "\in" "R"

av+bt "\in" T


Implying that a non- empty subset of v is a subspace if it is closed under linear combination


(b) Suppose u,v are in ker (f)


(1) f(u+v) = f(u)+f(v) = 0+0=0


(11) For a scalar, c

f(cv) =cf(v)=c.0=0


(iii)f(0)=f{(v)+(-v)}=f(v)+f(-v) =0+0=0


Therefore ker(f) is closed under zero, addition and multiplication


(c) Let y₁ and y2 "\in" im (f) and let r₁

,r2 be scalars


Consider r1y1+r2y2


There exist x1 and x2 "\in" v such

y1=f(x1) and y2=f(x2) and

that r₁f(x1)+r₂f(x2)= f(r₁x₁)+f(r2x2)


=f(r₁x1+r₂x2)


But f(r1x1+r2x2)"\in" Im (f)


Hence linear combination of y₁ and y2 is in Im (f)


Therefore Im (f) is a subspace of w


(d)


Let ker (f)={0}


Let f(x)=f(y) for x,y "\in" V


f(x)-f(y)=0

f(x-y)=0

"\therefore (x-y)\in" ker f


x-y=0

x=y


(e)


Let dim(Ker(f)) =k , dim(v)=n and n>k

Let Bo= {v1,...,vk} be a basis for ker (f)


Extending this basis to a full basis of V,

(n-k) vectors are added to the basis of the ker (f) to create a basis for V


Basis of V, Bv= {v1,...,vk,vk+1,...,vn}


The Image of f

Im [f]={f(v1),...,f(vk),f(v)k+1,...,f(vn)}


As v1,...,vk are in the kernel, f(v1),...,f(vk)=0

Therefore the basis for the Image of f is

{f(vk+1),...,f(vn)}



And it spans the range of f


Suppose there exists ak+1,…,an∈R ,not all zero such that

ak+1f(vk+1)+⋯+anf(vn)=0.


Using linearity of f, this is equivalent to f(ak+1vk+1+1+⋯+anvn)=0,

ak+1vk+1+⋯+anvn∈kerf


thus there exists b1,…,bk∈R


not all zero such that


ak+1vk+1+⋯+anvn=b1v1+⋯+bkvk


"\implies" All the constants are zero, which proves linear independence of


{f(vk+1),. . . . .f(vn)


"\therefore" dim ker f + dim Im [f] = dim v